OpenStudy (anonymous):

What would be the point-slope equation, using these two points: (-5, 3) and (4, -6)?

OpenStudy (anonymous):

Given that AC and AE are common external tangents of circle G and circle D, FE = 15, GF = 3, and AG = 5, which is the length of AC?

OpenStudy (anonymous):

slope formula is \[\frac{ y2-y1 }{ x2-x1 }\]

OpenStudy (anonymous):

so once you plug in ur numbers \[\frac{ -6-3 }{ 4-(-5) }\]

OpenStudy (anonymous):

After that we put the slope in the formula, \[y-y0= m(x-x0)\] ?

OpenStudy (anonymous):

do you mean onece you find the slop you have to put it into slope interscept form?

OpenStudy (anonymous):

Once I find the slope, I have to make a point-slope equation. I think that's the formula ^^

OpenStudy (anonymous):

the formula you have is point slope form that is right LOL sorry I got confused

OpenStudy (anonymous):

It's cool, I got it. Thanks! (:

OpenStudy (anonymous):

No Problem :)